Find the quantity of a 30% solution which must be mixed with a 15% solution to produce 600 kg of 25% solution
Answer by Alan3354(69443)

Find the quantity of a 30% solution which must be mixed with a 15% solution to produce 600 kg of 25% solution
-----------
t = amount of 30%
f = 15%
--------
t + f = 600
30t + 15f = 25*600
=========
etc